# Dominated Actions in Imperfect-Information Games

### v6

@@ Abstract
Dominance is a fundamental concept in game theory. In normal-form games dominated strategies can be identified in polynomial time. As a consequence, iterative removal of dominated strategies can be performed efficiently as a preprocessing step for reducing the size of a game before computing a Nash equilibrium. For imperfect-information games in extensive form, we could convert the game to normal form and then iteratively remove dominated strategies in the same way; however, this conversion may cause an exponential blowup in game size. In this paper we define and study the concept of dominated actions in imperfect-information games. Our main result is a polynomial-time algorithm for determining whether an action is dominated (strictly or weakly) by any mixed strategy in $n$-player games, which can be extended to an algorithm for iteratively removing dominated actions. This allows us to efficiently reduce the size of the game tree as a preprocessing step for Nash equilibrium computation. We explore the role of dominated actions empirically in “All In or Fold” No-Limit Texas Hold’em poker.
@@ 1 Introduction
In a normal-form game, (mixed) strategy $\sigma_{i}$ for player $i$ is strictly dominated if there exists another (mixed) strategy $\sigma^{\prime}_{i}$ such that $\sigma^{\prime}_{i}$ performs strictly better than $\sigma_{i}$ regardless of the strategies used by the opponent(s): formally, if $u_{i}(\sigma^{\prime}_{i},s_{-i})>u_{i}(\sigma_{i},s_{-i)}$ for all pure strategy profiles (vectors of pure strategies) $s_{-i}\in S_{-i}$ for the opposing agents. (Note that the requirement that this holds for all opposing pure strategy profiles is enough to ensure that it also holds for all mixed strategy profiles of the opponents as well.) Clearly it would be irrational for an agent to play a strictly dominated strategy, as another strategy will do strictly better regardless of the beliefs of what the opponents would play. Strategy $\sigma^{\prime}_{i}$ weakly dominates $\sigma$ if the inequality holds weakly (though strictly for at least one pure strategy profile): formally, $u_{i}(\sigma^{\prime},s_{-i})\geq u_{i}(\sigma_{i},s_{-i})$ for all $s_{-i}\in S_{-i}$ where the inequality is strict for at least one $s_{-i}$. (The condition that at least one inequality is strict is simply to rule out saying that a strategy is dominated by an identical strategy).
@@
There is a linear-time algorithm for determining whether a (mixed) strategy $\sigma_{i}$ is strictly dominated by any pure strategy for player $i$ [14]. This algorithm simply iterates over each pure strategy $s_{i}$ for player $i$ and tests whether it performs strictly better than $\sigma_{i}$ against each opposing pure strategy profile $s_{-i}$. This procedure has complexity $O(|A|)$, where $A=\times_{i}A_{i}$ is the set of joint pure strategy profiles, and so takes time linear in the size of the game. The procedure can also be easily adapted to produce an algorithm for determining whether a mixed strategy profile is weakly dominated by any pure strategy in linear time. Note that it is possible for a strategy to be dominated by a mixed strategy and not be dominated by any pure strategy [14]. In order to determine whether a (mixed) strategy is strictly dominated by a mixed strategy, while the above procedure does not work, it turns out that there exists a linear programming formulation that runs in polynomial time, and there also exists a linear programming formulation that determines whether a (mixed) strategy is weakly dominated by a mixed strategy [1]. Therefore, regardless of whether the strategies being tested as dominated or dominating are mixed or pure, it can be checked in polynomial time. [^1] However, most other computational questions related to dominance can be solved in polynomial time for strict but are NP-hard for weak dominance. For example, determining whether there exists some elimination path under which strategy $s_{i}$ is eliminated, given strategy subsets $S^{\prime}_{i}\subset S_{i}$ determining whether there exists a maximally reduced game where each player has the pure strategies $S^{\prime}_{i}$, and given constants $k_{i}$ determining whether there exists a maximally reduced game where each player has exactly $k_{i}$ pure strategies [14, 7].
Dominance is well-understood in normal-form games, where classical characterizations such as those of Pearce [12] describe exactly when a pure strategy is strictly or weakly dominated by a mixed strategy of the opponent. These characterizations, however, do not extend to extensive-form games. In imperfect-information settings, the dominance of an action depends on which opponent strategies reach the corresponding information set, a feature with no analogue in the normal-form framework. Consequently, an action in an extensive-form game may be locally dominated even when every pure or mixed strategy containing that action is undominated in the induced normal form. This distinction highlights the need for dominance notions and algorithms tailored specifically to extensive-form games.
@@ 2 Extensive-form games
While the normal form can be used to model simultaneous moves, another representation, called the extensive form, is generally preferred when modeling settings that have sequential moves. The extensive form can also model simultaneous moves, as well as chance events and imperfect information (i.e., situations where some information is available to only some of the agents and not to others). Extensive-form games consist primarily of a game tree; each non-terminal node has an associated player (possibly chance) that makes the decision at that node, and each terminal node has associated utilities for the players. Additionally, game states are partitioned into information sets, where the player whose turn it is to move cannot distinguish among the states in the same information set. Therefore, in any given information set, a player must choose actions with the same distribution at each state contained in the information set. If no player forgets information that they previously knew, we say that the game has perfect recall. A (behavioral) strategy for player $i,$ $\sigma_{i}\in\Sigma_{i},$ is a function that assigns a probability distribution over all actions at each information set belonging to $i$. In theory, every extensive-form game can be converted to an equivalent normal-form game; however, there is an exponential blowup in the size of the game representation, and therefore such a conversion is undesirable. Instead, new algorithms have been developed that operate on the extensive form representation directly. It turns out that the complexity of computing equilibria in extensive-form games is similar to that of normal-form games; a Nash equilibrium can be computed in polynomial time in two-player zero-sum games (with perfect recall) [8], while the problem is hard for two-player general-sum and multiplayer games. One algorithm for computing an equilibrium in two-player zero-sum extensive-form games with perfect recall is based on solving a linear programming formulation [8]. This formulation works by modeling each sequence of actions for each player as a variable, and is often called the sequence form LP algorithm. Note that while the number of pure strategies is exponential in the size of the game tree, the number of action sequences is linear. The method uses several matrices defined as follows. For player 1, the matrix $\mathbf{E}$ is defined where each row corresponds to an information set (including an initial row for the “empty” information set), and each column corresponds to an action sequence (including an initial row for the “empty” action sequence). In the first row of $\mathbf{E}$ the first element is 1 and all other elements are 0; subsequent rows have -1 for the entries corresponding to the action sequence leading to the root of the information set, and 1 for all actions that can be taken at the information set (and 0 otherwise). Thus $\mathbf{E}$ has dimension $c_{1}\times d_{1}$, where $c_{i}$ is the number of information sets for player $i$ and $d_{i}$ is the number of action sequences for player $i$. Matrix $\mathbf{F}$ is defined analogously for player 2. The vector $\mathbf{e}$ is defined to be a column vector of length $c_{1}$ with 1 in the first position and 0 in other entries, and vector $\mathbf{f}$ is defined with length $c_{2}$ analogously. The matrix $\mathbf{A}$ is defined with dimension $d_{1}\times d_{2}$ where entry $A_{ij}$ gives the payoff for player 1 when player 1 plays action sequence $i$ and player 2 plays action sequence $j$ multiplied by the probabilities of chance moves along the path of play. The matrix $\mathbf{B}$ of player 2’s expected payoffs is defined analogously. In zero-sum games $\mathbf{B}=-\mathbf{A}.$
@@
Given these matrices we can solve one of two linear programming problems to compute a Nash equilibrium in zero-sum extensive-form games [8]. In the first formulation the primal variables $\mathbf{x}$ correspond to player 1’s mixed strategy while the dual variables correspond to player 2’s strategy. In the second formulation, which is the dual problem of the first formulation, the primal decision variables $\mathbf{y}$ correspond to player 2’s strategy while the dual variables correspond to player 1’s strategy.
@@ 3 Dominated actions
The problem with these definitions is that they are too strong; it is still possible for player 1 to strictly prefer action $b_{i}$ to $a_{i}$ regardless of the strategy used by player 2 even if Candidate Definition 1 does not hold. This is illustrated in the following game depicted in Figure 1. In this game, chance makes an initial move taking each of two actions with probability $\frac{1}{2}.$ Then player 1 (red) selects one of two actions at a single information set. Player 2 (blue) then takes one of two actions after observing both the moves of chance and player 1. It is clear that action 2 for player 2 at their top information set and action 1 for player 2 at their bottom information set are both strictly dominated according to Candidate Definition 1. The smaller game obtained after removing these actions is depicted in Figure 2. In the smaller game, the expected utility of playing action 1 is 0.5(100)+0.5(100)=00.5(-100)+0.5(100)=0 and the expected utility of playing action 2 is 0.5(50)+0.5(50)=50.0.5(-50)+0.5(-50)=-50. Since player 2 does not take any actions, player 1 always achieves a strictly higher expected utility by playing action 1. However, action 2 is not strictly dominated according to Candidate Definition 1 because in one leaf node succeeding action 1 player 1 obtains payoff -100 which is lower than the payoff of -50 at leaf nodes following action 2.
@@
We next consider candidate definition given by Candidate Definition 3, where $u_{i}$ denotes the expected utility accounting for randomized moves of chance as well as potential randomization in the players’ strategies. This definition states that action $a_{i}$ for player $i$ at information set $I_{i}$ is strictly dominated (potentially by a probability distribution over other actions at the same information set), if there exists a strategy $\sigma^{-a_{i}}_{i}$ that never plays $a_{i}$ at $I_{i}$ that always has strictly higher expected utility than every strategy that plays $a_{i}$ at $I_{i}.$ Again this definition clearly provides a sufficient condition for $a_{i}$ to be dominated; however, the issue with this definition is that the strategies may potentially take actions early in the game tree that prevent the game from ever reaching information set $I_{i}.$ Consider the simple example game in Figure 3. Suppose we want to apply Candidate Definition 3 to determine whether action 2 is strictly dominated for player 2. Then $\sigma^{-a_{i}}_{i}$ is the strategy for player 2 that plays action 1 with probability 1, and $\sigma^{a_{i}}_{i}$ must be the strategy for player 2 that plays action 2 with probability 1. Now suppose that player 1 plays action 2 with probability 1, which we denote as $\sigma_{-i}$. Then clearly $u_{i}\left(\sigma^{-a_{i}}_{i},\sigma_{-i}\right)=u_{i}\left(\sigma^{a_{i}}_{i},\sigma_{-i}\right)=0,$ since both strategy profiles will result in reaching the bottom leaf node yielding payoff 0.
@@
Action $a_{i}$ for player $i$ is strictly dominated at information set $I_{i}$ if there exists a mixed strategy $\sigma^{-a_{i}}_{i}$ that plays action $a_{i}$ at $I_{i}$ with probability 0 such that for every mixed strategy $\sigma^{a_{i}}_{i}$ for player $i$ that plays action $a_{i}$ with probability 1 at $I_{i}$, $u_{i}\left(\sigma^{-a_{i}}_{i},\sigma_{-i}\right)>u_{i}\left(\sigma^{a_{i}}_{i},\sigma_{-i}\right)$ for all opposing strategies $\sigma_{-i}\in\Sigma_{-i}.$
@@
The problem with Candidate Definition 3 is that it allows the players to deviate from the path of play leading to the relevant information set $I_{i}.$ We address this limitation in our new definitions. Strictly-dominated actions are defined in Definition 1 and weakly-dominated actions are defined in Definition 2. In these definitions $\Sigma^{I_{i}}_{-i}$ denotes the set of mixed strategy profiles for the opponents that always take actions leading to information set $I_{i}$ when possible. Note that these definitions apply to games with any number of players. They also consider actions that are dominated by any mixed strategy (not necessarily just a pure action at $I_{i}$). It is easy to verify that these definitions address the issues that arose in the examples from Figure 1 and Figure 3. We can apply these definitions repeatedly in succession to perform iterative removal of dominated actions, as for the normal form.
@@
Action $a_{i}$ for player $i$ is strictly dominated at information set $I_{i}$ if there exists a mixed strategy $\sigma^{-a_{i}}_{i}$ that always plays to get to $I_{i}$ and plays action $a_{i}$ at $I_{i}$ with probability 0 such that for every mixed strategy $\sigma^{a_{i}}_{i}$ for player $i$ that always plays to get to $I_{i}$ and plays action $a_{i}$ with probability 1 at $I_{i}$, $u_{i}\left(\sigma^{-a_{i}}_{i},\sigma_{-i}\right)>u_{i}\left(\sigma^{a_{i}}_{i},\sigma_{-i}\right)$ for all opposing strategies $\sigma_{-i}\in\Sigma^{I_{i}}_{-i}.$
@@
Action $a_{i}$ for player $i$ is weakly dominated at information set $I_{i}$ if there exists a mixed strategy $\sigma^{-a_{i}}_{i}$ that always plays to get to $I_{i}$ and plays action $a_{i}$ at $I_{i}$ with probability 0 such that for every mixed strategy $\sigma^{a_{i}}_{i}$ for player $i$ that always plays to get to $I_{i}$ and plays action $a_{i}$ with probability 1 at $I_{i}$, $u_{i}\left(\sigma^{-a_{i}}_{i},\sigma_{-i}\right)\geq u_{i}\left(\sigma^{a_{i}}_{i},\sigma_{-i}\right)$ for all opposing strategies $\sigma_{-i}\in\Sigma^{I_{i}}_{-i}$, where the inequality is strict for at least one $\sigma_{-i}.$

### v7

@@ Abstract
Dominance is a fundamental concept in game theory. In normal-form games dominated strategies can be identified in polynomial time. As a consequence, iterative removal of dominated strategies can be performed efficiently as a preprocessing step for reducing the size of a game before computing a Nash equilibrium. For imperfect-information games in extensive form, we could convert the game to normal form and then iteratively remove dominated strategies in the same way; however, this conversion may cause an exponential blowup in game size. In this paper we define and study the concept of dominated actions in imperfect-information games. Our main result is a polynomial-time algorithm for determining whether an action is dominated (strictly or weakly) by any mixed strategy in two-player perfect-recall games with publicly observable actions, which can be extended to iteratively remove dominated actions. This allows us to efficiently reduce the size of the game tree as a preprocessing step for Nash equilibrium computation. We explore the role of dominated actions empirically in “All In or Fold” No-Limit Texas Hold’em poker.
@@ 1 Introduction
In a normal-form game, (mixed) strategy $\sigma_{i}$ for player $i$ is strictly dominated if there exists another (mixed) strategy $\sigma^{\prime}_{i}$ such that $\sigma^{\prime}_{i}$ performs strictly better than $\sigma_{i}$ regardless of the strategies used by the opponent(s): formally, if $u_{i}(\sigma^{\prime}_{i},s_{-i})>u_{i}(\sigma_{i},s_{-i)}$ for all pure strategy profiles (vectors of pure strategies) $s_{-i}\in S_{-i}$ for the opposing agents. (Note that the requirement that this holds for all opposing pure strategy profiles is enough to ensure that it also holds for all mixed strategy profiles of the opponents as well.) Clearly it would be irrational for an agent to play a strictly dominated strategy, as another strategy will do strictly better regardless of the beliefs of what the opponents would play. Strategy $\sigma^{\prime}_{i}$ weakly dominates $\sigma_{i}$ if the inequality holds weakly (though strictly for at least one pure strategy profile): formally, $u_{i}(\sigma^{\prime}_{i},s_{-i})\geq u_{i}(\sigma_{i},s_{-i})$ for all $s_{-i}\in S_{-i}$ where the inequality is strict for at least one $s_{-i}$. (The condition that at least one inequality is strict is simply to rule out saying that a strategy is dominated by an identical strategy).
@@
There is a linear-time algorithm for determining whether a (mixed) strategy $\sigma_{i}$ is strictly dominated by any pure strategy for player $i$ [14]. This algorithm simply iterates over each pure strategy $s_{i}$ for player $i$ and tests whether it performs strictly better than $\sigma_{i}$ against each opposing pure strategy profile $s_{-i}$. This procedure has complexity $O(|S|)$, where $S=\times_{i}S_{i}$ is the set of joint pure strategy profiles, and so takes time linear in the size of the game. The procedure can also be easily adapted to produce an algorithm for determining whether a mixed strategy profile is weakly dominated by any pure strategy in linear time. Note that it is possible for a strategy to be dominated by a mixed strategy and not be dominated by any pure strategy [14]. In order to determine whether a (mixed) strategy is strictly dominated by a mixed strategy, while the above procedure does not work, it turns out that there exists a linear programming formulation that runs in polynomial time, and there also exists a linear programming formulation that determines whether a (mixed) strategy is weakly dominated by a mixed strategy [1]. Therefore, regardless of whether the strategies being tested as dominated or dominating are mixed or pure, it can be checked in polynomial time. [^1] However, most other computational questions related to dominance can be solved in polynomial time for strict but are NP-hard for weak dominance. For example, determining whether there exists some elimination path under which strategy $s_{i}$ is eliminated, given strategy subsets $S^{\prime}_{i}\subset S_{i}$ determining whether there exists a maximally reduced game where each player has the pure strategies $S^{\prime}_{i}$, and given constants $k_{i}$ determining whether there exists a maximally reduced game where each player has exactly $k_{i}$ pure strategies [14, 7].
Dominance is well-understood in normal-form games, where classical characterizations such as those of Pearce [12] describe exactly when a pure strategy is strictly or weakly dominated by a mixed strategy of the opponent. These characterizations, however, do not extend to extensive-form games. In imperfect-information settings, the dominance of an action depends on which opponent strategies assign positive probability to histories reaching the corresponding information set, a feature with no analogue in the normal-form framework. Consequently, an action in an extensive-form game may be locally dominated even when every pure or mixed strategy containing that action is undominated in the induced normal form. This distinction highlights the need for dominance notions and algorithms tailored specifically to extensive-form games.
@@ 2 Extensive-form games
While the normal form can be used to model simultaneous moves, another representation, called the extensive form, is generally preferred when modeling settings that have sequential moves. The extensive form can also model simultaneous moves, as well as chance events and imperfect information (i.e., situations where some information is available to only some of the agents and not to others). Extensive-form games consist primarily of a game tree; each non-terminal node has an associated player (possibly chance) that makes the decision at that node, and each terminal node has associated utilities for the players. Additionally, game states are partitioned into information sets, where the player whose turn it is to move cannot distinguish among the states in the same information set. Therefore, in any given information set, a player must choose actions with the same distribution at each state contained in the information set. If no player forgets information that they previously knew, we say that the game has perfect recall. A game has publicly observable actions if, for every player, no information set contains nodes reached by different actions of another player. A (behavioral) strategy for player $i,$ $\sigma_{i}\in\Sigma_{i},$ is a function that assigns a probability distribution over all actions at each information set belonging to $i$. In theory, every extensive-form game can be converted to an equivalent normal-form game; however, there is an exponential blowup in the size of the game representation, and therefore such a conversion is undesirable. Instead, algorithms have been developed that operate on the extensive form representation directly. It turns out that the complexity of computing equilibria in extensive-form games is similar to that of normal-form games; a Nash equilibrium can be computed in polynomial time in two-player zero-sum games (with perfect recall) [8], while the problem is hard for two-player general-sum and multiplayer games.
One algorithm for computing an equilibrium in two-player zero-sum extensive-form games with perfect recall is based on solving a linear programming formulation [8]. This formulation works by modeling each sequence of actions for each player as a variable, and is often called the sequence form LP algorithm. Note that while the number of pure strategies is exponential in the size of the game tree, the number of action sequences is linear. The method uses several matrices defined as follows. For player 1, the matrix $\mathbf{E}$ is defined where each row corresponds to an information set (including an initial row for the “empty” information set), and each column corresponds to an action sequence (including an initial column for the “empty” action sequence). In the first row of $\mathbf{E}$ the first element is 1 and all other elements are 0; subsequent rows have -1 for the entries corresponding to the action sequence leading to the root of the information set, and 1 for all actions that can be taken at the information set (and 0 otherwise). Thus $\mathbf{E}$ has dimension $(c_{1}+1)\times(d_{1}+1)$, where $c_{i}$ is the number of information sets for player $i$ and $d_{i}$ is the number of action sequences for player $i$. Matrix $\mathbf{F}$ is defined analogously for player 2. The vector $\mathbf{e}$ is defined to be a column vector of length $c_{1}+1$ with 1 in the first position and 0 in other entries, and vector $\mathbf{f}$ is defined with length $c_{2}+1$ analogously. The matrix $\mathbf{A}$ is defined with dimension $(d_{1}+1)\times(d_{2}+1)$ where entry $A_{ij}$ gives the payoff for player 1 when player 1 plays action sequence $i$ and player 2 plays action sequence $j$ multiplied by the probabilities of chance moves along the path of play. The matrix $\mathbf{B}$ of player 2’s expected payoffs is defined analogously. In zero-sum games $\mathbf{B}=-\mathbf{A}.$
@@
Given these matrices we can solve one of two linear programming problems to compute a Nash equilibrium in zero-sum extensive-form games [8]. In the first formulation the primal variables $\mathbf{x}$ correspond to player 1’s strategy while the dual variables $\mathbf{q}$ correspond to player 2’s strategy. In the second formulation, which is the dual problem of the first formulation, the primal decision variables $\mathbf{y}$ correspond to player 2’s strategy while the dual variables $\mathbf{p}$ correspond to player 1’s strategy.
@@ 3 Dominated actions
The problem with these definitions is that they are too strong; it is still possible for player 1 to strictly prefer action $b_{i}$ to $a_{i}$ regardless of the strategy used by player 2 even if Candidate Definition 1 does not hold. This is illustrated in the following game depicted in Figure 1. In this game, chance makes an initial move taking each of two actions with probability $\frac{1}{2}.$ Then player 1 (red) selects one of two actions at a single information set. Player 2 (blue) then takes one of two actions after observing both the moves of chance and player 1. It is clear that action 2 for player 2 at their top information set and action 1 for player 2 at their bottom information set are both strictly dominated according to Candidate Definition 1. The smaller game obtained after removing these actions is depicted in Figure 2. In the smaller game, the expected utility of playing action 1 is 0.5(100)+0.5(100)=00.5(-100)+0.5(100)=0 and the expected utility of playing action 2 is 0.5(50)+0.5(50)=50.0.5(-50)+0.5(-50)=-50. Since player 2 does not take any actions, player 1 always achieves a strictly higher expected utility by playing action 1. However, action 2 is not strictly dominated according to Candidate Definition 1 because in one leaf node succeeding action 1 player 1 obtains payoff 100-100 which is lower than the payoff of 50-50 at leaf nodes following action 2.
@@
We next consider candidate definition given by Candidate Definition 3, where $u_{i}$ denotes the expected utility accounting for randomized moves of chance as well as potential randomization in the players’ strategies. This definition states that action $a_{i}$ for player $i$ at information set $I_{i}$ is strictly dominated (potentially by a probability distribution over other actions at the same information set), if there exists a strategy $\sigma^{-a_{i}}_{i}$ that never plays $a_{i}$ at $I_{i}$ that always has strictly higher expected utility than every strategy that plays $a_{i}$ at $I_{i}.$ Again this definition clearly provides a sufficient condition for $a_{i}$ to be dominated; however, the issue with this definition is that the strategies may potentially take actions early in the game tree that prevent the game from ever reaching information set $I_{i}.$ Consider the simple example game in Figure 3. Suppose we want to apply Candidate Definition 3 to determine whether action 2 is strictly dominated for player 2. Then $\sigma^{-a_{i}}_{i}$ is the strategy for player 2 that always plays action 1, and $\sigma^{a_{i}}_{i}$ must be the strategy for player 2 that plays action 2 with probability 1. Now suppose that player 1 plays action 2 with probability 1, which we denote as $\sigma_{-i}$. Then clearly $u_{i}\left(\sigma^{-a_{i}}_{i},\sigma_{-i}\right)=u_{i}\left(\sigma^{a_{i}}_{i},\sigma_{-i}\right)=0,$ since both strategy profiles will result in reaching the bottom leaf node yielding payoff 0.
@@
Action $a_{i}$ for player $i$ is strictly dominated at information set $I_{i}$ if there exists a behavioral strategy $\sigma^{-a_{i}}_{i}$ that plays action $a_{i}$ at $I_{i}$ with probability 0 such that for every behavioral strategy $\sigma^{a_{i}}_{i}$ for player $i$ that plays action $a_{i}$ with probability 1 at $I_{i}$, $u_{i}\left(\sigma^{-a_{i}}_{i},\sigma_{-i}\right)>u_{i}\left(\sigma^{a_{i}}_{i},\sigma_{-i}\right)$ for all opposing strategies $\sigma_{-i}\in\Sigma_{-i}.$
@@
The problem with Candidate Definition 3 is that it allows the players to deviate from the path of play leading to the relevant information set $I_{i}.$ We address this limitation in our new definitions. Strictly-dominated actions are defined in Definition 1 and weakly-dominated actions are defined in Definition 2. In these definitions $\Sigma^{I_{i}}_{-i}$ denotes the set of behavioral strategy profiles for the opponents that select only actions that do not prevent reaching $I_{i}$ whenever such actions exist. Note that these definitions apply to games with any number of players. They also consider actions that are dominated by any behavioral strategy (not necessarily just a pure action at $I_{i}$). It is easy to verify that these definitions address the issues that arose in the examples from Figure 1 and Figure 3. We can apply these definitions repeatedly in succession to perform iterative removal of dominated actions, as for the normal form.
@@
Action $a_{i}$ for player $i$ is strictly dominated at information set $I_{i}$ if there exists a behavioral strategy $\sigma^{-a_{i}}_{i}$ that always plays to get to $I_{i}$ and plays action $a_{i}$ at $I_{i}$ with probability 0 such that for every behavioral strategy $\sigma^{a_{i}}_{i}$ for player $i$ that always plays to get to $I_{i}$ and plays action $a_{i}$ with probability 1 at $I_{i}$, $u_{i}\left(\sigma^{-a_{i}}_{i},\sigma_{-i}\right)>u_{i}\left(\sigma^{a_{i}}_{i},\sigma_{-i}\right)$ for all opposing strategies $\sigma_{-i}\in\Sigma^{I_{i}}_{-i}.$
@@
Action $a_{i}$ for player $i$ is weakly dominated at information set $I_{i}$ if there exists a behavioral strategy $\sigma^{-a_{i}}_{i}$ that always plays to get to $I_{i}$ and plays action $a_{i}$ at $I_{i}$ with probability 0 such that for every behavioral strategy $\sigma^{a_{i}}_{i}$ for player $i$ that always plays to get to $I_{i}$ and plays action $a_{i}$ with probability 1 at $I_{i}$, $u_{i}\left(\sigma^{-a_{i}}_{i},\sigma_{-i}\right)\geq u_{i}\left(\sigma^{a_{i}}_{i},\sigma_{-i}\right)$ for all opposing strategies $\sigma_{-i}\in\Sigma^{I_{i}}_{-i}$, where the inequality is strict for at least one $\sigma_{-i}.$